OpenStudy (anonymous):

write as decimals 3/5= 6/10= 1 and 3/4=

OpenStudy (anonymous):

3/5=6/10=0.6 6/10=0.6 1 and 3/4 = 1.75

OpenStudy (saifoo.khan):

wow, he's right.

OpenStudy (anonymous):

it happens!

OpenStudy (saifoo.khan):

surely it does.

OpenStudy (anonymous):

GIVE ME YOUR BRAIN @_@

OpenStudy (saifoo.khan):

im emailing you, it's uploading.

OpenStudy (anonymous):

Really? =F

OpenStudy (anonymous):

Really? =F

OpenStudy (saifoo.khan):

yes!! =F

OpenStudy (anonymous):

take care of it numbers !

OpenStudy (saifoo.khan):

a photostate of it.

OpenStudy (saifoo.khan):

photostat*

OpenStudy (anonymous):

oh - not the original ?

OpenStudy (saifoo.khan):

then wht i will do? :O i think that will work fine as well.

OpenStudy (anonymous):

nice ;) thanks

OpenStudy (anonymous):

just post questions here and numbers will help you

OpenStudy (saifoo.khan):

hmm,

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

427,635,098.246 tell which digit is in the given place TENS: MILLIONS: HUNDREDTHS: TEN THOUSANDS:

OpenStudy (anonymous):

427,635,098.246 TENS: 9 MILLIONS: 7 HUNDREDTHS: 4 TEN THOUSANDS: 3

OpenStudy (anonymous):